OpenStudy (anonymous):

Can someone please check this? I fan and give medals! :) Which is an equation in slope-intercept form for the line passing through (5, 1) and (–4, 7). y=2/3x+13/3

OpenStudy (anonymous):

@Timelordguy

OpenStudy (anonymous):

I don't want to steal Timelordguy's thunder, but to check your answers, just plug in the values 5 and -4 as x into your conjecture (educated answer) and see if the values come out to y = 1 and 7 respectively.

OpenStudy (anonymous):

If you need assistance doing that, I am happy to help you.

OpenStudy (anonymous):

Yes actually I will need some help doing that. @IParke

OpenStudy (anonymous):

So, you believe your equation to be y=2/3x+13/3. To check your answer, plug in the coordinates you were supposed to have the equation fit. So, let's take the first coordinate your equation is supposed to satisfy: (5,1) This coordinate means that when x=5, y=1. Do you follow?

OpenStudy (anonymous):

Let's plug those values into your equation and see if the coordinate works.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y=2/3x+13/3 So, for (5,1), this becomes: (1) = (2/3)(5) + (13/3)

OpenStudy (anonymous):

Do you follow?

OpenStudy (anonymous):

yes I sorta get it

OpenStudy (anonymous):

So all I did was replace y with the 1 in (5,1) and the x with 5 in (5,1)

OpenStudy (anonymous):

Is this clear?

OpenStudy (anonymous):

Since (5,1) means x= 5 and y= 1.

OpenStudy (anonymous):

oh ok yes

OpenStudy (anonymous):

Good, now please evaluate the right side of your equation that I used substitution on: (1) = (2/3)(5) + (13/3)

OpenStudy (anonymous):

Tell me what the right side of the equation ends up being 1 = ?

OpenStudy (anonymous):

ok um well I just divide 2 by 3 then multiply by 5 and then add 13 divided by 3?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

7.66666666667

OpenStudy (anonymous):

Does 1 = 7.667?

OpenStudy (anonymous):

I am sorry to report, your answer is wrong.. BUT! I can give you a nice clue

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Make your 2/3's a negative, and try it again :)

OpenStudy (anonymous):

ok let me try again

OpenStudy (anonymous):

1=1

OpenStudy (anonymous):

I absolutely believe you know how to do this problem, you just made an error with the sign somewhere along the way :) You should find that 1=1.

OpenStudy (anonymous):

Excellent!

OpenStudy (anonymous):

great so it was correct?

OpenStudy (anonymous):

Now, for the final test, try using substitution with the other point: (–4, 7)

OpenStudy (anonymous):

So to recap, so far you've proven that the equation y=-2/3x+13/3 goes through (5,1). You now need to prove that it also goes through the other point, (-4, 7)

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

um yeah it kinda does

OpenStudy (anonymous):

Would you like me to walk you through proving that your equation goes through -4, 7?

OpenStudy (anonymous):

yes please your an amazing person!!! God bless you because my math skills are terrible!

OpenStudy (anonymous):

haha, God bless you too for trying to improve!

OpenStudy (anonymous):

So, we are going to use substitution again.

OpenStudy (anonymous):

haha thanks!

OpenStudy (anonymous):

So, please plug in the coordinate (-4,7) into the equation: y=-2/3x+13/3

OpenStudy (anonymous):

so would it be -4=-2/3(7)=13/3?

OpenStudy (anonymous):

Hmmm. Let us remember that the x-coordinate comes first in the coordinate pair :) (x = -4, y = 7)

OpenStudy (anonymous):

oh ok duh! so then its 7=-2/3(-4)+13/3

OpenStudy (anonymous):

@IParke

OpenStudy (anonymous):

Yes! Please evaluate the right side of the equation to see if 7=7!

OpenStudy (anonymous):

And if it does, then your equation is perfect :)

OpenStudy (anonymous):

Oh my gosh thank you so very much you ARE A LIFESAVER!!!!! :)

OpenStudy (anonymous):

Haha, I don't know about that, but I am so happy you're learning. If you are still confused, please let me know, I'll explain all of this to you in plain English.

OpenStudy (anonymous):

ok thanks I'll DEFINITELY need your help in the future!

OpenStudy (anonymous):

Good luck with your math endeavors!